OpenStudy (anonymous):

How do I solve this? 2x+10y=80 2x-5y=5 Help please!!

OpenStudy (anonymous):

I need to solve for x and y

OpenStudy (anonymous):

subtract the bottom equation from the top one then solve for y

OpenStudy (anonymous):

?

OpenStudy (anonymous):

So subtract 80 by 5

OpenStudy (anonymous):

2x + 10y = 80 -(2x + 5y = 5)

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

you would get 2x + 10y = 80 -(2x - 5y = 5) = 0x + 15y = 75

OpenStudy (anonymous):

So the answer would be 15

OpenStudy (anonymous):

Almost. You're left with 15y = 75. You now need to solve for what y is by dividing.

OpenStudy (anonymous):

If y is 15 I'll plug my answer back in for x

OpenStudy (anonymous):

So Y is 5. Will I plug that into 2x+10y=80?

OpenStudy (anonymous):

yes you can plug it in to either equation to find x

OpenStudy (anonymous):

So x is 7 right?

OpenStudy (anonymous):

just plug y into 2x -5y = 5 2x - 5(5) = 5 2x - 25 = 5 2x = 30 x = 15